OpenStudy (anonymous):

find a linear function f given f(5) = -1 and f(-5) = -3

OpenStudy (phi):

linear function means line they want f(x) = m x + b (that is the equation of a line) f(5) = -1 means (-1, 5) is a point on the line f(-5) = -3 tells you a second point on the line you have two points. Can you find the equation of a line through the two points ?

OpenStudy (gorv):

linear=> f(x)=ax+b f(5)=a*5+b=-1 -------(1) f(-5)=a*(-5)+b=-3 -5a+b=-3 b=-3+5a----(2) putting (2) in (1) 5a-3+5a=-1 10a=2 a=1/5 b=-2

OpenStudy (gorv):

@madams0032